OpenStudy (jekdidkdjrjrjjr):

HELP fan+medal for each question answered SEE PIC I need help with c,d,e........

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-boo

OpenStudy (jekdidkdjrjrjjr):

@queen.bee

OpenStudy (anonymous):

yes may i help you?

OpenStudy (jekdidkdjrjrjjr):

ya I need help

OpenStudy (jekdidkdjrjrjjr):

^

OpenStudy (bobo-i-bo):

For question c, substitue b=ik (where i is the imaginary number and k is an arbitary number) and see what happens! :)

OpenStudy (bobo-i-bo):

noope, i'm not doing it for you but i can help you :P

OpenStudy (jekdidkdjrjrjjr):

Like I'll make a new post for each question and give u a medal each time

OpenStudy (jekdidkdjrjrjjr):

Pls omfg I'm desperate

OpenStudy (jekdidkdjrjrjjr):

I'll like medal u rn

OpenStudy (bobo-i-bo):

nope :P

OpenStudy (jekdidkdjrjrjjr):

Whyyyyy

OpenStudy (jekdidkdjrjrjjr):

I'm so bad like I can't

OpenStudy (jekdidkdjrjrjjr):

I'll delete the question after so you won't get banned or something

OpenStudy (jekdidkdjrjrjjr):

I've been stuck on this forever like I need to finish dis

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-bo DUDE PLSZZZZKXDKKDKR

OpenStudy (bobo-i-bo):

Substitute b=ik! You haven't even tried!!

OpenStudy (jekdidkdjrjrjjr):

Sry but wtf is ik like wa

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-bo how do I do that??????!?!?

OpenStudy (bobo-i-bo):

Whenever there's a b, replace it with ik. Then solve! :)

OpenStudy (jekdidkdjrjrjjr):

ya I don't get it

OpenStudy (jekdidkdjrjrjjr):

@bobo-i-bo PLZ OMG

OpenStudy (bobo-i-bo):

\[a^2-b^2=a^2-(ik)^2\]

OpenStudy (jekdidkdjrjrjjr):

THATS THE ANSWER TYTYTYYYTYTYTT

OpenStudy (bobo-i-bo):

No that's not the answer ^_^

OpenStudy (jekdidkdjrjrjjr):

wtf

OpenStudy (bobo-i-bo):

That's the first step to getting the answer

OpenStudy (jekdidkdjrjrjjr):

ok, next step?

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-bo !?????

OpenStudy (jekdidkdjrjrjjr):

no offense

OpenStudy (bobo-i-bo):

Sorry, but i refuse to just you the answer '~'

OpenStudy (jekdidkdjrjrjjr):

Yeah but what's the next step?

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-bo

OpenStudy (bobo-i-bo):

\[a^2-(ik)^2=a^2-i^2k^2\]

OpenStudy (jekdidkdjrjrjjr):

That's it?

OpenStudy (jekdidkdjrjrjjr):

@bobo-I-bo ????YO